Project

General

Profile

NDOS Commissioning Meeting 2011/06/30

Attendance: Jon, Andrew, Gavin, Minerba, Mark Mathis, Mousumi, Athans, Gary, Martin, Raphael, Gareth, Dan, Mat, Denis

Agenda:

  1. Status of "Cool Running" (Jon)
  2. Status of NDOS DCS
  3. Status of APD repair
  4. AOB

Status of "Cool Running"

  • lost 5 cooled APDs late yesterday. One of these seems to have recovered after warming up. The other four are out of the DAQ readout.
  • Mat will be remove the latest failed FEBs tomorrow morning

Status of NDOS DCS

  • Athans:
    • IOC that Ron was working on is working with DSO scans. During scans, the IOCs are paused and renabled after a DSO scan.
    • Automatic start of cooling is not a feature. Must be manually activated.
    • Starting dcmApplication currently disables cooling
  • Jon: we will modify the script used by RC so that when shifters start a new run, they will have to click on the BeginRun button to get a run started. But before doing that, the shifter needs to make sure that cooling is enabled.
  • Athans: GUI is being developed; CSS "rules" are not picking up PVs properly, which is holding up making changes
  • Next DCS meeting is tomorrow, 9:30 am.

Status of APD repair

  • No news from CalTech
  • Mat: inspected some optical connectors of some APDs that had oil on them. One had a cracked optical connector. Attempted to "suck" oil out of the optical connector by hooking a vacuum; could not get any oil to come out.
  • Mat: we do have 5 APDs that cuold be installed, but we need to get install. procedures revamped
  • Jon: we should have a meeting to discuss the procedures, perhaps out at Lab 6 so others can watch the work in action.
  • Denis: one of the heatsinks would not cool, it looked like there was no water flow through the heatsink. Found water was blocked through the heatsink. Considering removing APD+heatsink and will send heatsink to IU.
  • Jon: since the APD is fine, let's hold off until
  • Denis: moved TECC lead to the outside of a special electronics box cover 3 different FEBs (same cover). On one, there was almost no noise. On other, no significant difference was seen.
  • Denis: 2 fuses on PDBs seem to be blown. 1 FEBs is getting no 3.3V power. The other fuse if for 24V power. Andrew will work with Martin on replacing these fuses.

DSO scan results

  • doc-dB 6719
  • looked at history of DSO scans of those APDs that went bad with cooling
  • label on last DSO scan for troublesome APDs should be "WARM" (they were not cooled during the last DSO scan)

AOB